# DIVERSE: A Dataset of YouTube Video Comment Stances with a Data Programming Model

### v2

@@ III-C6 Use of Analogous Stances
Finally, in the case of labeling the stance of the comment toward its respective video, we also used an analogous stance for a weak label. We used the stance of the comment toward the US Army. Generally speaking, if a comment was against the US Army, it would also be against a video posted by the US Army, and the use of complementary stances has been shown to improve performance in the stance detection task Sobhani et al. (2017); Zhang et al. (2023b). Thus, we used the final stance toward the US Army produced by weak supervision on the aforementioned weak labels as a weak label for the stance toward the video. It is important to note that this relationship does not work the other way; we found several examples where commentators would express a supportive stance toward the US Army but would be very against a particular video.
@@ III-E Human evaluation
After machine annotation, we had four individuals annotate how many? what percentage of total? comments. For each comment, the authors independently annotated the stance towards the US Army and the stance towards the video. The manual annotations were then analyzed for inter-annotator agreement, and the machine annotations were compared against the manual ones.
@@ IV Data Properties and Analysis
This section describes some exploratory data analysis performed on this dataset. The dataset consists of 858 videos published from 1 June 2010 through 5 October 2023 (final date of collection) on the US Army’s official YouTube Channel.
@@ IV-A Dataset Properties
Mean 16.77
Standard Deviation 32.67
@@
Mean 201.65
Standard Deviation 873.19
@@
25th Percentile 5
50th Percentile (Median) 27
75th Percentile 77
Maximum 16289
@@ IV-B Stance Label Properties
Interestingly, from the label counts we also observe both a relatively even distribution of labels across the stances but also the “neutral” stance being the minority label. In most other benchmark stance datasets, there is a strong label skew, particularly for datasets like Hossain et al. (2020) and Kochkina et al. (2018), the “neutral” stance tends to be the majority stance label. Since this is a first-of-its-kind dataset for both the targets as well as the data domain (i.e., YouTube) it’s not clear if the targets tend to inspire stronger stance-taking behavior in comments or if the domain of YouTube causes stronger stance-taking behavior. Regardless, upon manual inspection of some of the labels, we do find that these stance labels are accurate and that there is a lot of strong stance-taking within the comments.
@@
against 63676 (34.87%) 67405 (36.92%)
neutral 49153 (26.91%) 49310 (27.01%)
supports 60183 (32.99%) 56297 (30.88%)
@@
Turning to the distribution of stance labels over time, we find that each of the three classes of stance generally trend together over time. Figure 4 displays the distribution of the different classes of stance toward the US Army for all of the comments. The different stances toward the videos follow a very similar trend. We further observe that the two prominent periods of comments do differ somewhat in their compositions of stance labels. In the first peak of comments, centering in September of 2022, we see more “against” stance labels than “supports” or “neutral”. Whereas in the second spike of comments around May-June of 2023, we see more “supports” comments than “against” or “neutral”. Thus, overall, there is a moderate balance among the stances expressed in the comments toward the videos, but this balance does shift over time and between videos.
@@
To better understand these differences in stances, we looked at the videos with the most positive, neutral and negative receptions. Among the videos with more than 100 comments, we found the video, “Honoring Our Fallen Heroes - Taps” had the highest percentage of supportive comments at 76.5% of the comments supporting the video. On the opposite end was the video “What does the Army of the future look like?” which features a short interview with former Sergeant Major Grinston on the future of the US Army. The video had 66.9% of comments that were against it. And, the most neutrally received video was “Army #PopQuiz - What training is this?” which received 66.5% of its comments as being neutral toward the video, many of them answering the question posed in the video.
@@ IV-C Comparison with human evaluation
The inter-annotator agreement was Iain, results here. Comparing the machine-based annotations with the human-based annotations, we get an accuracy of how many. elaborate on accuracy
IV-D Leave-One-Out Stance Detection
To further validate the stance labels produced by the proposed weak supervision method, we performed a leave-one-out, out-of-domain experiment referenced from (Ng and Carley, 2022). We trained a model on the stance labels and then evaluated that model on the other benchmark datasets. In our test, we trained a cased BERT model [^9] https://huggingface.co/bert-base-cased (Devlin et al., 2018) on only the stance labels from the DISMISS dataset and then evaluated the model on all of the benchmark datasets. The model was trained for 3 epochs using a learning rate of $41e-$ 4 with linear decay, a batch size of 32, and was trained to classify text into three different labels. The text that was input to the model was given as:
“target: {target of stance} [SEP] comment: {comment}”
In this test, if the proposed data labels are capturing the stances correctly, then we should see similar performance on the benchmark datasets when predicting stance, as was seen in Ng and Carley (2022). Table IV has the out-of-domain test results. The results are evaluated using the Macro F1 score as a suitable evaluation method due to the almost equal distribution of stance labels.
On average, the model trained on stances only from this dataset achieves a macro F1 score of 0.3043. This is similar to the average accuracy of 0.303 in a past work that performed leave-one-out stance detection using multiple BERT-based classifiers Ng and Carley (2022). This therefore suggests that the stance labels proposed in this paper are valid stance labels.
Dataset Macro F1 Score
semeval2016 0.3517
election2016 0.3439
covid-lies 0.3020
phemerumors 0.2295
wtwt 0.3018
srq 0.2971
Average 0.3043
TABLE IV: Macro F1 Scores for Model Applied to Different Datasets
@@ V Discussion
Public opinion towards a country’s army is crucial, as it signals the level of trust in the army and the willingness of citizens to join the forces. A positive opinion facilitates army recruitment and support, especially in times of crisis. Social media engagement plays a pivotal role in shaping this opinion. This paper introduces a dataset that supports such understanding and associated analyses, comprising comments on the US Army’s official YouTube videos annotated for their stance towards the Army and the video content itself.
@@
Finally, our Leave-One-Out experiment produces a standard macro F1 score of around 0.34, a score that is similar to past work. This suggests that the stance labels produced by the proposed methodology are indeed valid. This result also highlights findings from past work that trained stance detection models do not necessarily generalize to other data domains (e.g., YouTube versus Twitter, or targets of videos versus political figures, etc.). Thus, we suggest training a stance classification model using multiple benchmark datasets so the model can better recognize the nuances of the stances and targets involved. The low F1 score also indicates the difficulty of the stance classification task, whereby language models do struggle to differentiate the nuances in opinions towards a target entity, despite having helper annotations like hate and sentiment to provide indications of stances. This, therefore, opens up avenues for exploration of improving the stance classification methods, especially on out-of-domain tasks, to provide a more accurate label towards an entity and therefore more accurate downstream insights.
@@
The dataset proposed in this work has several potential uses. First, this dataset presents a versatile resource for research on the analysis of stance in YouTube video comments. Beyond its primary application in training and evaluating new stance detection techniques and models, the dataset introduces distinctive traits not present in current benchmarks. Notably, it originates from YouTube, offering a departure from the prevalent X-based datasets. Furthermore, the dataset features multiple stance targets and exhibits a substantial degree of stance-taking behavior and controversial content (i.e., a high proportion of “supports” and “against” stances). The dataset provides sequential information about comments, allowing investigations into the dynamics of stance with respect to the stance expressed in a previous comment.

### v3

@@ III-C6 Use of Analogous Stances
Finally, in the case of labeling the stance of the comment toward its respective video, we also used an analogous stance for a weak label; we used the stance of the comment toward the U.S. Army as a weak label for the stance toward the video. Generally speaking, if a comment was against the U.S. Army, it would also be against a video posted by the U.S. Army, and the use of complementary stances has been shown to improve performance in the stance detection task Sobhani et al. (2017); Zhang et al. (2023b). Thus, we used the final stance toward the U.S. Army produced by weak supervision on the aforementioned weak labels as a weak label for the stance toward the video. It is important to note that this relationship does not work the other way; we found several examples where commentators would express a supportive stance toward the U.S. Army but would be very against a particular video.
@@ III-E Human evaluation
In order to assess the quality of the machine-produced labels, we needed to validate these labels by humans. Since it is impractical to label the entire dataset by humans, we randomly selected 1,000 comments which were labeled by four annotators. We then assessed the label agreement among the annotators and combined the human-labels by majority vote. Using these labels, we assessed the quality of the machine-produced labels.
@@ IV Data Properties and Analysis
This section describes some exploratory data analysis performed on this dataset. The dataset consists of 1082 videos published from 1 June 2010 through 1 April 2024 on the US Army’s official YouTube Channel. From these videos, there are a total of 215,509 comments.
@@ IV-A Dataset Properties
Mean 16.66
Standard Deviation 33.13
@@
Mean 199.17
Standard Deviation 883.41
@@
25th Percentile 7
50th Percentile (Median) 37
75th Percentile 86
Maximum 18999
@@ IV-B Stance Label Properties
From the labeling we can see some label skew. First, we observe a label skew toward the “neutral” label class, which is characteristic of all other stance-labeled datasets (e.g., Hossain et al. (2020) and Kochkina et al. (2018)). We do also, however, observe a slight label skew toward “against” stances versus “supportive” stances. Since these are comments meant to advocate for a brand (i.e., the U.S. Army), the skew in valenced stances against the U.S. Army and its videos is a cause for concern. Finally, we also observe that through the use of the weak labels, we were unable to label to a couple of thousand data points (around $1\%$ of the data) for each stance.
@@
abstain 2373 (1.10%) 1564 (0.73%)
against 49817 (23.12%) 63222 (29.34%)
neutral 122028 (56.61%) 102479 (47.56%)
supports 41291 (19.16%) 48244 (22.37%)
@@
Turning to the distribution of stance labels over time, we find that each of the three classes of stance generally trend together over time. Figure 4 displays the distribution of the different classes of stance toward the US Army for all of the comments. The different stances toward the videos follow a very similar trend. We further observe that the two prominent periods of comments do differ somewhat in their compositions of stance labels. In the first peak of comments, centering in September of 2022, we see more “against” stance labels than “supports”. Whereas in the second spike of comments around May-June of 2023, we see more “supports” comments than “against”. Thus, overall, there is a moderate balance among the stances expressed in the comments toward the videos, but this balance does shift over time and between videos.
@@
To better understand these differences in stances, we looked at the videos with the most positive, neutral and negative receptions. Among the videos with more than 100 comments, we found the video, “Honoring Our Fallen Heroes - Taps” had the highest percentage of supportive comments at 61.5% of the comments supporting the video. On the opposite end was the video “Sergeant Major of the Army Grinston addresses the report of the Fort Hood Independent Review” which features a short interview with former Sergeant Major Grinston on the investigation at fort Hood following the murder of the Venessa Guillen. The video had 55.9% of comments that were against it. And, the most neutrally received video was “Army #PopQuiz - What training is this?” which received 81.6% of its comments as being neutral toward the video, many of them answering the question posed in the video.
@@ IV-C Comparison with human evaluation
Finally, we analyzed how well machine labeling, using state–of–the–art techniques, of stance, performed. To do this, we created a validation dataset as previously described and had four human labelers label that data set. For the stance toward the U.S. Army, we had an average Cohen’s Kappa McHugh (2012) across all of the human labelers of $0.863$ and for the stance toward the videos of $0.797$. Thus, we were able to obtain high agreement between the labelers on stance of the comments (note that we also released this labeled validation dataset).
Compared to the output from the weak supervision process, we obtained moderately successful machine labels. For the stance toward the U.S. Army of the comments, the accuracy and Micro F1-scores were $0.68$ and for the stance toward the videos of the comments, the accuracy and Micro F1-scores were $0.639$. When we similarly compared the results of each of the LLMs in the labeling task (see Table IV), we do see an improvement of the weak supervision-produced labels over the best-performing LLMs labels, especially for the stance toward the videos.
Labeling Model U.S. Army Video
mistral zero-shot with comments 66.90% 56.50%
llama-3 zero-shot with comments 41.40% 50.90%
mistral zero-shot 64.40% 55.10%
llama-3 zero-shot 41.20% 40.20%
ul2 zero-shot with comments 53.00% 45.00%
ul2 zero-shot 49.00% 45.80%
mistral adapter 49.70% 26.50%
small model 52.50% 32.30%
TABLE IV: Micro-F1 scores for U.S. Army and Video stances, by different models and prompts, on the validation dataset.
From the analysis of the LLM labelers, we notice that there are distinct differences in performance between models and prompting schemes, with the Mistral LLM Jiang et al. (2023) and the prompt scheme incorporating replies in the comments generally performing better.
To further understand these results, we plotted the confusion matrix between the human-derived and machine-derived labels in Figure 6. From these matrices, we can observe that for valenced stances, the machine-label stances generally match those from human-annotation; much of loss of accuracy in the machine-labeled stances comes from mis-labeling the neutrally-stanced comments. Upon inspection of some these examples, it seems there are a number of highly-positive sounding comments that are unrelated to the video or the Army as well as some comments where sarcasm or coded language (e.g., ‘goy’, ‘zog’ were missed by the LLMs.
Figure 6: Confusion matrices between the machine-derived labels and the human-derived labels for the validation set.
@@ V Discussion
Public opinion towards a country’s military is crucial, as it signals the level of trust in the government writ large and the willingness of citizens to join the military. A positive opinion facilitates military recruitment and support, especially in times of crisis. Social media engagement plays a pivotal role in shaping this opinion. This paper introduces a dataset that supports such understanding and associated analyses, comprising comments on the US Army’s official YouTube videos annotated for their stance towards the Army and the video itself.
@@
In this work we combined the existing state-of-the-art methods for automated labeling of a dataset on the difficult task of stance labeling. We found that these approaches were able to produce reasonable labels, relative to a human-labeled validation dataset, but also that there is still room for improvement in the development of automated labeling techniques, especially for challenging labeling tasks, like stance. This, therefore, opens up avenues for exploration of improving the stance classification methods, especially on out-of-domain tasks, to provide a more accurate label towards an entity and therefore more accurate downstream insights.
@@
The dataset proposed in this work has several potential uses. First, this dataset presents a versatile resource for research on the analysis of stance in YouTube video comments. Beyond its primary application in training and evaluating new stance detection techniques and models, the dataset introduces distinctive traits not present in current benchmarks. Notably, it originates from YouTube, offering a departure from the prevalent X or Twitter-based datasets. Furthermore, the dataset features multiple stance targets and exhibits a substantial degree of stance-taking behavior and controversial content (i.e., a high proportion of “supports” and “against” stances). The dataset provides sequential information about comments, allowing investigations into the dynamics of stance with respect to the stance expressed in a previous comment.